OpenStudy (anonymous):

Given the curve how do I parametrize the tangent line at (2,1,8)?

OpenStudy (loser66):

Take derivative of the curve. Plug it into the formula of the tangent line at the point.

OpenStudy (anonymous):

I know the derivative would be <1/(2radt),-4/t^2,t> but what would the formula be @Loser66

OpenStudy (loser66):

tangent line : x =x0 + at y = y0 + bt z = z0+ ct, where <a, b, c> is derivative of the curve, <x0,y0,z0> is the point

OpenStudy (anonymous):

should I be plugging something in for t? @Loser66

OpenStudy (loser66):

or you can use this formula, it works well also ;) a(x-x0)+b(y-y0) +c(z-z0)=0

OpenStudy (anonymous):

so would it be (1/2radt)(x-2)-4(y-1)/t^2+t(z-8)? @Loser66

OpenStudy (loser66):

I think so. :)

OpenStudy (loser66):

=0

OpenStudy (anonymous):

alrite thank you :)